Two blocks P (9 kg) and Q (2 kg) connected by a light string are being pulled along a horizontal frictionless surface by a horizontal force F=78 N. What is tension force along the string?

N.
Answer #1

- PON 10 F-T= ma 78-T = ga - T= maa = T = 2a - ② Addig & 78 = lla y a= 78% from @ T=2a = 2x78 ~ 14.18 N Ang